OpenStudy (anonymous):

find the length of the missing side. leave your answer in simplest radical form a. 2 sqrt 41 b. 164 c. 6 d. 2

OpenStudy (anonymous):

@csscgirl

OpenStudy (nurali):

According to Pythagorean, a² + b² = c² a=8 and c=10 c²-a²= b² 10²-8²=b² b²=36 b²=6

OpenStudy (anonymous):

6

OpenStudy (anonymous):

well, c